Differentiate the following functions with respect to x:
(i) ln(1 + sin2 x) (ii) x²
1
Expert's answer
2020-06-29T18:51:01-0400

(i)

Let y = ln(1+sin2x)

Also let z = 1+sin2x

So y = ln(z)

Differentiating with respect to z

"\frac{dy}{dz}" = "\frac{1}{z}"

Again z = 1+sin2x

So "\frac{dz}{dx}" = 2cos2x as "\frac{d(1)}{dx} = 0" and ", \frac{d}{dx}[sin(mx)] = mcos(mx)"

Now "\frac{dy}{dx} = \frac{dy}{dz} \frac{dz}{dx}" = "\frac{1}{z}" (2cos 2x)

=> "\frac{dy}{dx} = \frac{2cos2x}{1+sin2x}"

You can further simplify as follows

"\frac{dy}{dx} = \frac{2cos2x(1-sin2x)}{1-sin^22x}"

=> "\frac{dy}{dx} = \frac{2cos2x(1-sin2x)}{cos^22x}"

=> "\frac{dy}{dx} = \frac{2(1-sin2x)}{cos2x}"

=> "\frac{dy}{dx} = 2(sec2x-tan2x)"

Therefore

"\frac{d}{dx}(ln(1+sin2x)"

= "2(sec2x-tan2x)"

(ii)

Let y = x²

We know that "\frac{d}{dx}(x^n) = nx^{n-1}"

So "\frac{dy}{dx} = 2x^{2-1}"

=> "\frac{dy}{dx} = 2x"

So "\frac{d}{dx}(x^2) = 2x"